### v1

@@ 4.3 Ablation Study and Component Analysis
To isolate the specific contribution of each preprocessing module within the proposed dual pipeline framework, we conducted a comprehensive ablation study. This analysis systematically evaluates how the incremental addition of normalization, class balancing, feature selection, and dimensionality reduction influences model performance. Table 10 and Table 11 present the performance trajectories for the representative models of Pipeline 1 (K Nearest Neighbors) and Pipeline 2 (Extra Trees), respectively.
Table 10: Stepwise ablation study of the K Nearest Neighbors classifier within Pipeline 1 evaluating the impact of statistical feature engineering
@@
Table 11: Stepwise ablation study of the Extra Trees classifier within Pipeline 2 evaluating the impact of wrapper based selection and autoencoders
@@ 4.3.1 Efficacy of Hybrid Resampling
The SMOTETomek hybrid resampling technique was employed to enhance recall rates for the minority classes (Insomnia and Sleep Apnea) by synthetically generating minority samples while simultaneously cleaning class boundaries. The impact of this module varied significantly across model architectures. For distance based algorithms like K Nearest Neighbors and boosting frameworks like LightGBM, the integration of SMOTETomek yielded substantial gains. Specifically, Table 10 shows that combining SMOTETomek with Mutual Information feature selection elevated the K Nearest Neighbors accuracy from 96.0% to 98.667% and the F1 Score to 97.850%. Conversely, linear models such as Logistic Regression did not benefit from this oversampling strategy, observing a slight degradation in performance metrics. This suggests that while resampling is critical for defining local decision boundaries in non parametric models, it provides limited utility for models relying on global linear separation.
@@ 4.3.2 Impact of Data Normalization
The necessity of feature scaling proved to be highly model dependent. The Multi Layer Perceptron classifier exhibited a critical dependency on normalization, improving from a baseline accuracy of 61.333% to 96.000% after the application of RobustScaler. Similarly, the K Nearest Neighbors model improved from 92.0% to 94.667% accuracy upon scaling, as the algorithm requires uniform feature magnitude to calculate Euclidean distances effectively. In contrast, tree based ensembles such as Random Forest and Extra Trees demonstrated inherent robustness to feature scaling, maintaining consistent accuracy regardless of the normalization technique applied (RobustScaler or MinMaxScaler). These findings confirm that while normalization is a mandatory preprocessing step for neural and distance based classifiers, it is redundant for decision tree ensembles.
@@
We compared the efficacy of filter based Mutual Information against the wrapper based Boruta algorithm. Mutual Information proved particularly effective for models sensitive to noise, such as K Nearest Neighbors and AdaBoost. As indicated in Table 10, the inclusion of Mutual Information features enabled the K Nearest Neighbors model to achieve its peak performance by filtering out redundant dimensions that distort distance calculations. Although the Extra Trees model in Pipeline 2 achieved the highest global accuracy, we emphasize the utility of the K Nearest Neighbors configuration in Pipeline 1. This model was selected for optimization not only for its high accuracy but also because the Mutual Information selection process is computationally less expensive than the iterative Boruta method, a trade off we analyze further in the computational efficiency section. In contrast, Boruta was primarily effective for the Extra Trees classifier (Table 11), where it successfully identified all relevant features during the oversampling process, boosting the F1 Score to 97.840%.
@@
The application of dimensionality reduction techniques yielded mixed results across the classifiers. Linear Discriminant Analysis in Pipeline 1 successfully projected the feature space for the K Nearest Neighbors and Random Forest models, maintaining a high accuracy of 96.0%. However, the projection into lower dimensions caused a loss of variance that negatively impacted Logistic Regression and Gradient Boosting. Similarly, the unsupervised Autoencoder in Pipeline 2 generally resulted in suboptimal performance for tree based models compared to the original feature space. However, the Autoencoder successfully denoised the data for the Multi Layer Perceptron and K Nearest Neighbors classifiers, allowing them to achieve competitive accuracy in a compressed feature space. This suggests that while aggressive dimensionality reduction can streamline the feature space for neural networks, it may eliminate critical decision boundary information required by ensemble tree classifiers.
@@ 4.4 Computational Efficiency Analysis
A significant disparity in training efficiency was observed across the classifiers. In Pipeline 1, the K Nearest Neighbors model configured with Mutual Information feature selection demonstrated exceptional efficiency, requiring only 1.04 seconds for training while achieving the highest accuracy of 98.67%. This contrasts sharply with the XGBoost model in the same pipeline, which required over 901 seconds to converge due to the computational intensity of gradient boosting on the expanded feature set. Consequently, K Nearest Neighbors represents the optimal balance of speed and precision for the statistical engineering pipeline. Within the wrapper based framework of Pipeline 2, the Extra Trees classifier (184.90 seconds) proved significantly faster than XGBoost (1923.86 seconds) while matching its superior accuracy. Although simple linear models like Logistic Regression offered the lowest latency (12.59 seconds), their inability to capture non linear dependencies resulted in lower diagnostic sensitivity compared to the ensemble methods. Therefore, despite the higher computational cost relative to linear baselines, the Extra Trees architecture is justified by its significant performance gains and acceptable training duration.
For deployment in real time clinical monitoring systems, the testing or prediction time is paramount. As indicated in the tables, all optimized models achieved inference times well below 400 milliseconds. The Multi Layer Perceptron and Logistic Regression models provided near instantaneous predictions (approximately 2 to 3 milliseconds). Even the complex ensemble models maintained rapid inference speeds suitable for immediate feedback, ensuring that the proposed framework is viable for continuous patient monitoring applications.
@@ 4.5 Overfitting Assessment and Generalization Behavior
Ensemble Regularization: The best performing model, Extra Trees (Extremely Randomized Trees), is less prone to overfitting than a single decision tree and often more robust than highly flexible neural models on small tabular datasets. Unlike standard Random Forests that search for an optimal split at each node, Extra Trees selects split thresholds at random. This added randomness slightly increases bias but reduces variance, which helps prevent the model from fitting spurious patterns in a limited dataset. In our experiments, training and validation performance remained close across the cross validation folds.
Strict Separation in Preprocessing: To prevent data leakage, all preprocessing steps, including scaling and feature selection (Boruta and Mutual Information), were fit only on the training portion of each cross validation fold and then applied to its corresponding validation portion. This ensures that no information from validation data influences preprocessing decisions. In addition, Boruta filters out features that do not consistently outperform shadow features, which further limits the model capacity to fit noise.
Efficacy of Stratified Cross Validation: We employed stratified eight fold cross validation so that each fold preserves the original class proportions. The stability of performance across folds (standard deviation $1.5\%$) suggests that the results are not driven by a single favorable split and are consistent across different subsets of the data.
@@ 4.6 Statistical Significance Verification
To validate that the performance improvements achieved by our proposed frameworks are not artifacts of random variance in the data splitting process, we performed the Wilcoxon Signed Rank Test on the fold wise accuracy scores from the Stratified 8 Fold Cross Validation. This non parametric statistical hypothesis test is preferred over the Paired t test as it makes no assumption regarding the normality of the accuracy distribution.We conducted two distinct pairwise comparisons to verify the efficacy of both processing pipelines. First, we compared the baseline K Nearest Neighbors model (92.00% mean accuracy) against the optimized Pipeline 1 configuration utilizing Mutual Information and SMOTETomek (98.67% mean accuracy). Second, we compared the baseline Extra Trees model (96.00% mean accuracy) against the optimized Pipeline 2 configuration utilizing Boruta and SMOTETomek (98.67% mean accuracy).
The analysis yielded a p value of 0.00391 ($W=36.0$) for both comparisons. Since this p value is significantly lower than the standard significance level of $\alpha=0.05$, we reject the null hypothesis in both instances. This result statistically confirms that the proposed dual pipeline framework provides a genuine and robust improvement in diagnostic accuracy over standard baseline classifiers, verifying that the gains are systematic rather than circumstantial.
Table 14: Comparative analysis of the proposed framework against recent state of the art studies in sleep disorder classification
@@
Han et al.(2023)[54] Sleep Clinic of Samsung Medical Center K-means Clustering 91%
@@ 4.7 Benchmarking against State-of-the-Art
To contextualize the diagnostic capability of the proposed Dual Pipeline Framework, we compared our results against recent state of the art studies employing similar methodologies or datasets. Table 14 presents a comparative analysis involving twelve related works, focusing on studies utilizing the Sleep Health and Lifestyle Dataset. The proposed Extra Trees classifier, optimized via Pipeline 2 (Boruta and SMOTETomek), achieved a classification accuracy of 98.667%, establishing a new benchmark for this dataset. Notably, this level of performance was not unique to the tree based ensemble; the K Nearest Neighbors, XGBoost, and LightGBM models from Pipeline 1 also attained this identical peak accuracy of 98.667%. We selected the Extra Trees classifier for the primary comparison in Table 14 to illustrate the efficacy of the wrapper based feature selection pipeline. Our framework demonstrates a quantifiable improvement over the most recent benchmarks established in 2025. Specifically, the model outperforms the Gradient Boosting approach by Rahman et al. (2025), which achieved 97.33%, and the Ensemble Model by Monowar et al. (2025), which reported 96.88% accuracy. Furthermore, the proposed method significantly surpasses deep learning implementations on this tabular data, such as the Artificial Neural Network approach by Alshammari et al. (2024), which yielded 92.92%. This comparison validates that the rigorous application of hybrid resampling and targeted feature engineering in our dual pipeline architecture provides a superior solution for automated sleep disorder screening.
@@ 6 Conclusion
Our empirical results confirmed that the rigorous integration of hybrid SMOTETomek resampling is essential for mitigating algorithmic bias against minority classes such as Insomnia and Sleep Apnea. The framework achieved a state of the art classification accuracy of 98.67% using Extra Trees and K Nearest Neighbors classifiers, a performance improvement statistically validated by the Wilcoxon Signed Rank Test. Beyond diagnostic precision, the proposed model demonstrates profound practical utility for large scale computerized screening. With inference latencies maintained below 400 milliseconds, the system meets the stringent requirements for real time clinical monitoring, facilitating immediate risk stratification and early intervention prior to the onset of severe comorbidities.
Despite these promising results, the study is limited by the relatively small sample size and reliance on self reported lifestyle metrics. Future research will focus on validating this framework against larger multi center clinical datasets to ensure demographic universality. Furthermore, we aim to adapt this architecture for integration with wearable medical devices, creating a hybrid modality that combines continuous physiological time series with static lifestyle profiling. Ultimately, deploying this lightweight and high sensitivity model into mobile health ecosystems could democratize access to sleep disorder diagnostics, enabling identification outside of conventional clinical settings and significantly reducing the long term burden on healthcare infrastructure.
@@
  • Conflict of Interest: The authors have no relevant financial or non financial interests to disclose.

### v2

@@ 4.3 Ablation Study and Component Analysis
To isolate the specific contribution of each preprocessing module within the proposed dual pipeline framework, we conducted a comprehensive ablation study. This analysis systematically evaluates how the incremental addition of normalization, class balancing, feature selection, and dimensionality reduction influences model performance. Table 10 and Table 11 present the performance trajectories for the representative models of Pipeline 1 (K-Nearest Neighbors) and Pipeline 2 (Extra Trees), respectively.
Table 10: Stepwise ablation study of the K Nearest Neighbors classifier within Pipeline 1, evaluating the impact of statistical feature engineering
@@
Table 11: Stepwise ablation study of the Extra Trees classifier within Pipeline 2, evaluating the impact of wrapper-based selection and autoencoders
@@ 4.3.1 Efficacy of Hybrid Resampling
The SMOTETomek hybrid resampling technique was employed to enhance recall rates for the minority classes (Insomnia and Sleep Apnea) by synthetically generating minority samples while simultaneously cleaning class boundaries. The impact of this module varied significantly across model architectures. For distance-based algorithms like K Nearest Neighbors and boosting frameworks like LightGBM, the integration of SMOTETomek yielded substantial gains. Specifically, Table 10 shows that combining SMOTETomek with Mutual Information feature selection elevated the K Nearest Neighbors accuracy from 96.0% to 98.667% and the F1 Score to 97.850%. Conversely, linear models such as Logistic Regression did not benefit from this oversampling strategy, observing a slight degradation in performance metrics. This suggests that while resampling is critical for defining local decision boundaries in non-parametric models, it provides limited utility for models relying on global linear separation.
@@ 4.3.2 Impact of Data Normalization
The necessity of feature scaling proved to be highly model-dependent. The Multi-Layer Perceptron classifier exhibited a critical dependency on normalization, improving from a baseline accuracy of 61.333% to 96.000% after the application of RobustScaler. Similarly, the K Nearest Neighbors model improved from 92.0% to 94.667% accuracy upon scaling, as the algorithm requires uniform feature magnitude to calculate Euclidean distances effectively. In contrast, tree-based ensembles such as Random Forest and Extra Trees demonstrated inherent robustness to feature scaling, maintaining consistent accuracy regardless of the normalization technique applied (RobustScaler or MinMaxScaler). These findings confirm that while normalization is a mandatory preprocessing step for neural and distance-based classifiers, it is redundant for decision tree ensembles.
@@
We compared the efficacy of filter-based Mutual Information against the wrapper-based Boruta algorithm. Mutual Information proved particularly effective for models sensitive to noise, such as K-Nearest Neighbors and AdaBoost. As indicated in Table 10, the inclusion of Mutual Information features enabled the K-Nearest Neighbors model to achieve its peak performance by filtering out redundant dimensions that distort distance calculations. Although the Extra Trees model in Pipeline 2 achieved the highest global accuracy, we emphasize the utility of the K-Nearest Neighbors configuration in Pipeline 1. This model was selected for optimization not only for its high accuracy but also because the Mutual Information selection process is computationally less expensive than the iterative Boruta method, a trade-off we analyze further in the computational efficiency section. In contrast, Boruta was primarily effective for the Extra Trees classifier (Table 11), where it successfully identified all relevant features during the oversampling process, boosting the F1 Score to 97.840%.
@@
The application of dimensionality reduction techniques yielded mixed results across the classifiers. Linear Discriminant Analysis in Pipeline 1 successfully projected the feature space for the K-Nearest Neighbors and Random Forest models, maintaining a high accuracy of 96.0%. However, the projection into lower dimensions caused a loss of variance that negatively impacted Logistic Regression and Gradient Boosting. Similarly, the unsupervised Autoencoder in Pipeline 2 generally resulted in suboptimal performance for tree-based models compared to the original feature space. However, the Autoencoder successfully denoised the data for the Multi Layer Perceptron and K-Nearest Neighbors classifiers, allowing them to achieve competitive accuracy in a compressed feature space. This suggests that while aggressive dimensionality reduction can streamline the feature space for neural networks, it may eliminate critical decision boundary information required by ensemble tree classifiers.
@@ 4.4 Computational Efficiency Analysis
A significant disparity in training efficiency was observed across the classifiers. In Pipeline 1, the K-Nearest Neighbors model configured with Mutual Information feature selection demonstrated exceptional efficiency, requiring only 1.04 seconds for training while achieving the highest accuracy of 98.67%. This contrasts sharply with the XGBoost model in the same pipeline, which required over 901 seconds to converge due to the computational intensity of gradient boosting on the expanded feature set. Consequently, K-Nearest Neighbors represents the optimal balance of speed and precision for the statistical engineering pipeline. Within the wrapper-based framework of Pipeline 2, the Extra Trees classifier (184.90 seconds) proved significantly faster than XGBoost (1923.86 seconds) while matching its superior accuracy. Although simple linear models like Logistic Regression offered the lowest latency (12.59 seconds), their inability to capture nonlinear dependencies resulted in lower diagnostic sensitivity compared to the ensemble methods. Therefore, despite the higher computational cost relative to linear baselines, the Extra Trees architecture is justified by its significant performance gains and acceptable training duration.
For deployment in real-time clinical monitoring systems, the testing or prediction time is paramount. As indicated in the tables, all optimized models achieved inference times well below 400 milliseconds. The Multi-Layer Perceptron and Logistic Regression models provided near instantaneous predictions (approximately 2 to 3 milliseconds). Even the complex ensemble models maintained rapid inference speeds suitable for immediate feedback, ensuring that the proposed framework is viable for continuous patient monitoring applications.
@@ 4.5 Overfitting Assessment and Generalization Behavior
Ensemble Regularization: The best performing model, Extra Trees (Extremely Randomized Trees), is less prone to overfitting than a single decision tree and often more robust than highly flexible neural models on small tabular datasets. Unlike standard Random Forests that search for an optimal split at each node, Extra Trees selects split thresholds at random. This added randomness slightly increases bias but reduces variance, which helps prevent the model from fitting spurious patterns in a limited dataset. In our experiments, training and validation performance remained close across the cross-validation folds.
Strict Separation in Preprocessing: To prevent data leakage, all preprocessing steps, including scaling and feature selection (Boruta and Mutual Information), were fit only on the training portion of each cross-validation fold and then applied to its corresponding validation portion. This ensures that no information from validation data influences preprocessing decisions. In addition, Boruta filters out features that do not consistently outperform shadow features, which further limits the model’s capacity to fit noise.
Efficacy of Stratified Cross Validation: We employed stratified eight-fold cross-validation so that each fold preserves the original class proportions. The stability of performance across folds (standard deviation $1.5\%$) suggests that the results are not driven by a single favorable split and are consistent across different subsets of the data.
@@ 4.6 Statistical Significance Verification
To validate that the performance improvements achieved by our proposed frameworks are not artifacts of random variance in the data splitting process, we performed the Wilcoxon Signed Rank Test on the fold-wise accuracy scores from the Stratified 8 Fold Cross Validation. This non-parametric statistical hypothesis test is preferred over the Paired t test as it makes no assumption regarding the normality of the accuracy distribution. We conducted two distinct pairwise comparisons to verify the efficacy of both processing pipelines. First, we compared the baseline K Nearest Neighbors model (92.00% mean accuracy) against the optimized Pipeline 1 configuration utilizing Mutual Information and SMOTETomek (98.67% mean accuracy). Second, we compared the baseline Extra Trees model (96.00% mean accuracy) against the optimized Pipeline 2 configuration utilizing Boruta and SMOTETomek (98.67% mean accuracy).
The analysis yielded a p-value of 0.00391 ($W=36.0$) for both comparisons. Since this p-value is significantly lower than the standard significance level of $\alpha=0.05$, we reject the null hypothesis in both instances. This result statistically confirms that the proposed dual pipeline framework provides a genuine and robust improvement in diagnostic accuracy over standard baseline classifiers, verifying that the gains are systematic rather than circumstantial.
Table 14: Comparative analysis of the proposed framework against recent state-of-the-art studies in sleep disorder classification
@@ 4.7 Benchmarking against State-of-the-Art
To contextualize the diagnostic capability of the proposed Dual Pipeline Framework, we compared our results against recent state-of-the-art studies employing similar methodologies or datasets. Table 14 presents a comparative analysis involving eleven related works, focusing on studies utilizing the Sleep Health and Lifestyle Dataset. The proposed Extra Trees classifier, optimized via Pipeline 2 (Boruta and SMOTETomek), achieved a classification accuracy of 98.667%, establishing a new benchmark for this dataset. Notably, this level of performance was not unique to the tree-based ensemble; the K Nearest Neighbors, XGBoost, and LightGBM models from Pipeline 1 also attained this identical peak accuracy of 98.667%. We selected the Extra Trees classifier for the primary comparison in Table 14 to illustrate the efficacy of the wrapper-based feature selection pipeline. Our framework demonstrates a quantifiable improvement over the most recent benchmarks established in 2025. Specifically, the model outperforms the Gradient Boosting approach by Rahman et al. (2025), which achieved 97.33%, and the Ensemble Model by Monowar et al. (2025), which reported 96.88% accuracy. Furthermore, the proposed method significantly surpasses deep learning implementations on this tabular data, such as the Artificial Neural Network approach by Alshammari et al. (2024), which yielded 92.92%. This comparison validates that the rigorous application of hybrid resampling and targeted feature engineering in our dual pipeline architecture provides a superior solution for automated sleep disorder screening.
@@ 6 Conclusion
Our empirical results confirmed that the rigorous integration of hybrid SMOTETomek resampling is essential for mitigating algorithmic bias against minority classes such as Insomnia and Sleep Apnea. The framework achieved a state of the art classification accuracy of 98.67% using Extra Trees and K Nearest Neighbors classifiers, a performance improvement statistically validated by the Wilcoxon Signed Rank Test. Beyond diagnostic precision, the proposed model demonstrates profound practical utility for large-scale computerized screening. With inference latencies maintained below 400 milliseconds, the system meets the stringent requirements for real-time clinical monitoring, facilitating immediate risk stratification and early intervention prior to the onset of severe comorbidities.
Despite these promising results, the study is limited by the relatively small sample size and reliance on self-reported lifestyle metrics. Future research will focus on validating this framework against larger multi center clinical datasets to ensure demographic universality. Furthermore, we aim to adapt this architecture for integration with wearable medical devices, creating a hybrid modality that combines continuous physiological time series with static lifestyle profiling. Ultimately, deploying this lightweight and high-sensitivity model into mobile health ecosystems could democratize access to sleep disorder diagnostics, enabling identification outside of conventional clinical settings and significantly reducing the long term burden on healthcare infrastructure.
@@
  • Conflict of Interest: The authors have no relevant financial or non-financial interests to disclose.